Solve this application problem using a system of equations: Ron and Wanda are ticket sellers for their class play. Ron sells student tickets for $2.00 each, and Kathy sells adult tickets for $4.50 each. If their total revenue for 369 tickets is $1175.50, then how many tickets did Ron sell?
Answer by mananth(16946)

Number of student tickets sold ----x
Number of adult tickets ------------y
x+y = 369
multiply by 4.5
4.5x+4.5y = 1660.50
2x+ 4.5y = 1175.50
subtract
2.5x =585
x=234